Until recently, Marco's Pizza sold a large pizza for $20 and they sell 250 large pizzas per week. They just raised the price to $28 and as a result, they only sell 150 pizzas per week. Using the midpoint rule, find the price elasticity of demand for Marco's Pizza's large pizzas. Is the demand for Marco's elastic or inelastic? Did their total revenue increase or decrease with the change in price?
1
Expert's answer
2018-06-05T08:48:08-0400
P1 = $20, Q1 = 250 pizzas per week. P2 = $28, Q2 = 150 pizzas per week. The price elasticity of demand for Marco's Pizza's large pizzas is:
Ed = (150 - 250)/(28 - 20)*(28 + 20)/(150 + 250) = -100/8*48/400 = 1.5, so the demand for Marco's is elastic, that's why their total revenue will decrease with the change in price.